OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3). y = three halvesx - 1 y = three halvesx + 3 y = -two thirdsx - 1 y = -two thirdsx + 3

OpenStudy (texaschic101):

y = 3/2x + 4 the slope of this line is 3/2. A perpendicular line will have the negative reciprocal slope. All that means is " flip " the slope and change the sign. So the slope we need is -2/3 (see how I flipped the slope and changed the sign) Now we will use y = mx + b slope(m) = -2/3 (-6,3) x = -6 and y = 3 now we sub 3 = -2/3(-6) + b 3 = 4 + b 3 - 4 = b -1 = b so our perpendicular line will be : y = -2/3x - 1

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thank you so much!

OpenStudy (texaschic101):

no problem...glad to help

OpenStudy (anonymous):

btw how do you give a medal

OpenStudy (texaschic101):

click on best response by the person you want to give a medal to

OpenStudy (anonymous):

thx again @texaschic101

OpenStudy (texaschic101):

anytime :)